OpenStudy (anonymous):

differentiate w.r.t x. 1+ 3sec^2 x the answer should be 3sec^2 y tan y

OpenStudy (anonymous):

you're wondering how to do it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know the quotient rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use it

OpenStudy (anonymous):

I assume you're supposed to show how you do it instead of just magically coming up with the answer

OpenStudy (anonymous):

so there is no substitution involved

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what will be my denominator

OpenStudy (amistre64):

isnt there a rule for sec(x), and a power rule for ^n and a chain rule as well?

OpenStudy (anonymous):

you can do it using 3/sin^2

OpenStudy (anonymous):

it should be 3/cos^2 x

OpenStudy (amistre64):

sec = 1/cos ... not seeing the 1/sin^2

OpenStudy (anonymous):

sorry, my mistake

OpenStudy (anonymous):

ok give me 2 minutes

OpenStudy (amistre64):

we learn by mistakes ... and ive made plenty :)

OpenStudy (anonymous):

lol am making one now lol

OpenStudy (amistre64):

the quotient rule will work ... i just think its more hassle than its worth for this problem.

OpenStudy (anonymous):

numerator i have -6cosx

OpenStudy (anonymous):

denominator cos^4 x

OpenStudy (anonymous):

i have made a slight mistake in the question i don't know if it significant the equation is 1+3 sec^2 y

OpenStudy (anonymous):

so implicit differentiation

OpenStudy (amistre64):

implicit is just the same stuff ... but a chain rule pops out a y'

OpenStudy (amistre64):

the 1 is useless since its a constant; and constants derive to 0 3 sec^2 (y) might be seen better as: (sec y) ^2 run a power rule: 3*2 (sec y), and out pops the insides 6 (sec y) * (sec y)' the derivatie of sec is sectan 6 sec y * sec y tan y and out pops the insides .. 6 sec y * sec y tan y * y'

OpenStudy (anonymous):

the answer is given as 3 sec^2 y tan y

OpenStudy (amistre64):

im not all that concerned with "what the answer given" is .... the rules of derivatives dont change regardless of "what the given answer" is.

OpenStudy (amistre64):

make sure youve posted the correct information, thats all i have to go off of

OpenStudy (anonymous):

ok let me post the whole question

OpenStudy (anonymous):

if sin y= 2 sin x show that (a) (dy/dx)^2= 1+ 3sec^2 x (i have solved this part as you need to differentiate and then substitute by differentiating (a) w.r.t.x show that d^y/dx^2=3sec2^y tan y

OpenStudy (anonymous):

error on a should be (dy/dx)^2=1+ 3 sec^2 y

OpenStudy (amistre64):

sin y= 2 sin x cos(y) y' = 2 cos(x) y' = 2 cos(x)/cos(y) y' = 2 cos(arcsin (sin(y)/2) )/cos(y) that looks fun

OpenStudy (amistre64):

how did you go about that?

OpenStudy (anonymous):

firstly we do the substitution for sin y=sqrt 1-cos^2 y

OpenStudy (anonymous):

likewise for 2 sin x=sqrt 1-cos^ 2 x

OpenStudy (anonymous):

at the end you should get the expression 4 cos^2 x= 3+cos^2 x

OpenStudy (amistre64):

\[sin(y)=u\] \[cos(y)~y'=u'\] \[y'=\frac{u'}{cos(y)}\] \[y'=\frac{u'}{\sqrt{1-(u)^2}}\] \[y'=\frac{2~cos(x)}{\sqrt{1-(2~sin(x))^2}}\] thats ringing a few bells :)

OpenStudy (anonymous):

you the differentiate the original function implicitly and you end up with dy/dx=4cos^x/cos^y then substitute

OpenStudy (anonymous):

finally you should end up with 3/cos^2 y + 1 which gives you your answer to part a

OpenStudy (amistre64):

that looks fine to me :) and then theres this part: d^y/dx^2 = 3sec2^y tan y , this is odd notation

OpenStudy (anonymous):

sorry am a bit exhausted d^2 y/dx^2=3sec^2 y tan y

OpenStudy (amistre64):

so the second derivative; and do they want you to use 1+3sec^2 y for it? or the results of the first derivativing?

OpenStudy (anonymous):

yes you are just asked to find the second derivative of the function

OpenStudy (amistre64):

\[y'=\frac{2~cos(x)}{\sqrt{1-(2~sin(x))^2}}\] \[y'=2~cos(x)(1-4~sin^2(x))^{(-1/2)}\] \[y''=-2~sin(x)(1-4~sin^2(x))^{(-1/2)}+8cos^2(x)sin(x)~(1-4~sin^2(x))^{(-3/2)}\] looks to be accurate

OpenStudy (amistre64):

\[y''=\frac{-2~sin(x)\sqrt{1-4~sin^2(x)}+8cos^2(x)sin(x)}{(1-4~sin^2(x))~\sqrt{1- 4~sin^2(x)}}\] lol, thats looks to be fun to process

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=2nd+derivative+arcsin%282sin%28x%29%29 might help to see it simplified

OpenStudy (anonymous):

thanks